# 1st Stanislav Krikunov Skoltech Moscow, Russia

### v1

@@ I Introduction
Meanwhile, nonlinear devices can be either passive or active. IMD products of passive devices are called Passive Intermodulation (PIM). Among PIM sources are weak mechanical connections in the TX chain, kinks and sharp edges in conductors, duplexer filters or ferrite fillers, switches, metal oxide layers covering conductors and junctions, and dirt in connectors. Passive nonlinearity is mainly caused by nonlinear conductive and magnetic properties of devices inside the transceiver chain (internal PIM) or outside the system in the near-field or transition antenna array region (external PIM, air-induced PIM) due to metal fences or billboards near the antenna array [4].
PIM represents one of the major interference problems in modern radio systems for service providers and equipment suppliers. PIM interference results in decreased coverage of BS cells, a decrease in the sensitivity of RX uplink (UL) signals, or possibly the complete inoperable transmission link.
@@ III-B External PIM compensation in 16T16R MIMO
The result for a large-scale MIMO system with 16 transmitting and receiving antennas is shown. Each transceiver chain in the system is represented by a set of half-wavelength dipoles grouped into unit cells of 4 antennas. The signal from each channel is divided equally among all antennas within a unit cell. The distance between the antennas is equal to half of the wavelength (w.r.t. the central frequency of the TX signal band). Even channels have vertical polarization; odd channels have horizontal polarization (crossed dipoles). The antenna layout is shown in Fig. 3. A dashed line limits unit cells.
III-B1 Single point source compensation.
The compensation ability of the algorithm for scenario 1 is shown in Fig. 4. External PIM is cancelled to the noise floor after full algorithm convergence. The algorithm convergence curve is shown in Fig. 5.
III-B2 Multiple point source compensation.
@@
III-C Observed effects
III-C1 Near-field PIM power variation
The location of the PIM source significantly affects the received power distribution between the channels. This is especially noticeable in the near-field zone of the antenna array since the wave here is spherical (Fig. 7) and the electric field varies a lot. As the source moves away from the antenna array, the wave becomes flat, and such a power variation is no longer noticeable (Fig. 8). Polarization effects can also be noticed since the signal power at even antennas differs from the signal of odd antennas.
III-C2 Identification of the number of PIM sources
The technique described in paper [13] can be used to calculate the number of independent PIM sources. The authors apply Eigenvalue Decomposition (EVD) to obtain eigenvalues of the RX signal covariance matrix. The number of eigenvalues exceeding the noise power level corresponds to the number of PIM sources spatially distinguishable from the receiver side. Generally, the number of eigenvalues exceeding the noise floor eigenvalues does not correspond to the real number of objects generating PIM. For example, 2 closely located sources may generate similar signals by level and structure. Therefore, in terms of eigenvalues, they can be considered the same source of external PIM.
Several cases of the number of external PIM sources are demonstrated in Fig. 9: 1, 2 and 4 sources. The eigenvalues of the RX covariance matrix for each case are sorted. The external PIM average power is artificially set to the same value by scaling in each case.
Figure 7: RX PIM power variation at the distance of 0.1 m from the antenna array (high power variation).
Figure 8: RX PIM power variation at the distance of 1 m from the antenna array (low power variation).
Figure 9: Eigenvalues of the covariance matrix for external PIM generated by 1 / 2 / 4 point sources.

### v2

@@ I Introduction
Meanwhile, nonlinear devices can be either passive or active. IMD products of passive devices are called Passive Intermodulation (PIM). Among PIM sources are weak mechanical connections in the TX chain, kinks and sharp edges in conductors, duplexer filters or ferrite fillers, switches, metal oxide layers covering conductors and junctions, and dirt in connectors. Passive nonlinearity is mainly caused by nonlinear conductive and magnetic properties of devices inside the transceiver chain (internal PIM) or outside the system in the near-field [12] or transition antenna array region (external PIM, air-induced PIM) due to metal fences or billboards near the antenna array [4].
PIM represents one of the major interference problems [13, 14, 15] in modern radio systems for service providers and equipment suppliers. PIM interference results in decreased coverage of BS cells, a decrease in the sensitivity of RX uplink (UL) signals, or possibly the complete inoperable transmission link.
@@ III-B External PIM compensation in 16T16R MIMO
The result for a large-scale MIMO system with 16 transmitting and receiving antennas is shown. Each transceiver chain in the system is represented by a set of half-wavelength dipoles grouped into unit cells of 4 antennas. The signal from each channel is divided equally among all antennas within a unit cell. The distance between the antennas is equal to half of the wavelength (w.r.t. the central frequency of the TX signal band). Even channels have vertical polarization; odd channels have horizontal polarization (crossed dipoles). The antenna layout is shown in Fig. 3 where numbering means antenna number which is not related to the number of TX chains directly. A dashed line limits unit cells.
III-B1 Single point source compensation
The compensation ability of the algorithm for scenario 1 is shown in Fig. 4. External PIM is cancelled to the noise floor after full algorithm convergence (Fig. 5).
III-B2 Multiple point source compensation
@@
III-C Near-field PIM power variation
The location of the PIM source significantly affects the received power distribution between the channels. This is especially noticeable in the near-field zone of the antenna array since the wave here is spherical and the electric field varies a lot. As the source moves away from the antenna array, the wave becomes flat, and such a power variation is no longer noticeable (Fig. 7). Polarization effects can also be noticed since the signal power at even antennas differs from the signal of odd antennas.
Figure 7: RX PIM power variation at the distance of 0.1 m from the antenna array (left, high power variation) and 1m (right, low power variation).